# SonicSieve: Bringing Directional Speech Extraction to Smartphones Using Acoustic Microstructures

### v2

@@ 4.2. Subjective user evaluation of audio quality
While objective metrics like SI-SDR provide valuable quantitative assessment, they cannot fully capture the perceptual quality of speech enhancement systems in real-world scenarios. To evaluate the actual listening experience, we conduct a user study to complement our objective measurements. We randomly select 12 generated audio samples from our test set collected from Room A. The selected audio contains one target speech and 1-2 interfering sources (undesired speech or other sound). Each sample is processed by three systems: the 2-channel microphone array (baseline), 5-channel microphone array, and our proposed SonicSieve. 20 participants (14 males and 6 females) with an average age of 24 years were recruited for the listening test. The evaluation protocol consisted of three steps: (1) listening to the noisy audio mixture, (2) listening to the clean target speaker audio as reference, and (3) listening to the three processed outputs presented in random order to avoid position bias. Participants were asked to rate each processed audio sample on two criteria using 5-point scales. We note that two separate questions (Recommendation, 2003) are used to decouple the assessment of noise removal from the preservation of target speech quality, as aggressive noise suppression can often introduce undesirable artifacts that a single question would fail to capture.
@@
Figure 14. User study measuring listening experience via mean opinion scores: (a) audibility of interfering speakers and background noise, (b) overall listening quality.
[…]
This figure displays results from a user study measuring listening experience through mean opinion scores on a 1-5 scale (higher is better), presented in two side-by-side bar charts with error bars: (a) Left panel ”Noise Suppression Performance” compares three systems: 2-mic array: Mean score of approximately 2.5 with error bars ranging from about 1.3 to 3.7. 5-mic array: Mean score of approximately 3.3 with error bars ranging from about 2.1 to 4.6. SonicSieve: Mean score of approximately 4.2 with error bars ranging from about 2.1 to 4.4. (b) Right panel ”Overall Experience” compares the same three systems: 2-mic array: Mean score of approximately 2.3 with error bars ranging from about 1.3 to 3.5. 5-mic array: Mean score of approximately 3.1 with error bars ranging from about 1.9 to 4.4. SonicSieve: Mean score of approximately 3.1 with error bars ranging from about 2.1 to 4.5. Panel (a) specifically measures audibility of interfering speakers and background noise (noise suppression capability), while panel (b) measures overall listening quality. The data suggests that the SonicSieve system performs comparably to or slightly better than the 5-mic array across both metrics, while both significantly outperform the 2-mic array. The overlapping error bars indicate some variability in user opinions, but the general trend shows preference for the SonicSieve and 5-mic systems.
As shown in Fig. 14, both SonicSieve and the 5-channel microphone array significantly outperform the 2-channel microphone baseline in subjective evaluations. For noise suppression performance (Fig. 14(a)), SonicSieve achieved a Mean Opinion Score (MOS) of 3.25 $\pm$ 1.16, comparable to the 5-microphone array’s 3.33 $\pm$ 1.21, despite using fewer microphones. Statistical analysis using a Student’s t-test indicated that the difference was not statistically significant (t = 0.77, p = 0.44; $\alpha$ = 0.05). The overall listening experience (Fig. 14(b)) shows that SonicSieve outperforms the 5-microphone array (3.35 $\pm$ 1.23 vs. 3.10 $\pm$ 1.22, t=-2.19, p=0.03; $\alpha$ = 0.05). This statistical significance confirms that participants consistently preferred SonicSieve’s performance in real-world listening scenarios.
@@ 5. LIMITATIONS AND DISCUSSION
  • Generalization across microphone frequency responses. To enable generalization across different smartphones microphones, device-specific calibration can be performed by measuring the frequency response of the built-in and in-line microphones, and deriving equalization weights to normalize for their effects. This can allow our system to robustly perform directional speech extraction across a range of devices. A complementary approach is the use of established data-driven approaches such as training on diverse device recordings (Li et al., 2017b; Ardila et al., 2019) and targeted data augmentation techniques (Morocutti et al., 2023) which may enable zero-shot generalization without calibration for unseen devices.
@@
Source separation within a sector. Our system currently performs spatial filtering along predefined directional sectors and produces a mixture of all sounds it. This can be undesirable when multiple speakers are close to each other and fall within the same sector. Future extensions of this work could address this limitation in two ways: (1) A low-latency speech separation framework (Gu et al., 2021, 2022; Yuan et al., 2024) can be applied after directional speech extraction, using permutation-invariant training to separate multiple speakers. (2) More fine-grained separation can be performed in the spatial domain by building on techniques for region-customizable sound extraction (Gu and Luo, 2024) to spatially isolate the target speaker.
@@
A limitation of the current in-line microphone setup is the close proximity of the mic to the earphones, resulting in insufficient wire length to comfortably reach a user’s ears when the smartphone is placed on a table. This can be addressed in two ways. First, a custom but low-cost headphone could be designed with the in-line mic positioned farther from the speakers, allowing the phone to remain on a surface while maintaining sufficient wire length. Alternatively, wireless headphones could be used for playback by integrating the acoustic microstructure onto one of the outward-facing microphones typically used for noise cancellation. This approach is likely possible as modern headphones often include dual outward-facing microphones for noise cancellation, and low-power GPUs are becoming integrated into wearable devices (Tambe et al., 2022; gpu, 2025).

### v3

@@ 4.2. Subjective user evaluation of audio quality
While objective metrics like SI-SDR provide valuable quantitative assessment, they cannot fully capture the perceptual quality of speech enhancement systems in real-world scenarios. To evaluate the actual listening experience, we conduct a user study to complement our objective measurements. This study was approved by our Institutional Review Board (STUDY2025_00000146). All studies complied with relevant ethical regulations.
We randomly select 25 generated audio samples from our test set collected from three different rooms (Room A-C). The selected audio contains one target speech and 1-2 interfering sources (undesired speech or other sound). Each sample is processed by three systems: the 2-channel microphone array (baseline), 5-channel microphone array, and our proposed SonicSieve. 44 participants in total (31 males and 13 females) with an average age of 25 years were recruited for the listening test. The evaluation protocol consisted of three steps: (1) listening to the noisy audio mixture, (2) listening to the clean target speaker audio as reference, and (3) listening to the three processed outputs presented in random order to avoid position bias. Participants were asked to rate each processed audio sample on two criteria using 5-point scales. We note that two separate questions (Recommendation, 2003) are used to decouple the assessment of noise removal from the preservation of target speech quality, as aggressive noise suppression can often introduce undesirable artifacts that a single question would fail to capture.
@@
As shown in Fig. 14, both SonicSieve and the 5-channel microphone array significantly outperform the 2-channel microphone baseline in subjective evaluations. Across the 25 samples drawn from three rooms, for noise suppression performance (Fig. 14(a)), SonicSieve achieved a Mean Opinion Score (MOS) of 3.25 $\pm$ 1.20, compared to the 5-microphone array’s 3.46 $\pm$ 1.24 and the 2-channel baseline’s 2.48 $\pm$ 1.17. For overall listening experience (Fig. 14(c)), SonicSieve achieved 3.39 $\pm$ 1.20, comparable to the 5-microphone array (3.35 $\pm$ 1.19) and higher than the 2-channel baseline (2.52 $\pm$ 1.13). Fig. 14(c) further breaks down overall experience by room, showing consistent gains over the 2-channel baseline across Room A–C, with SonicSieve remaining comparable to the 5-channel array in each environment.
4.3. User Interface Evaluation
Figure 15. User interface evaluation setup for directional speech extraction, where the participants are asked to select one or both speakers on the smartphone UI with different configurations in different environments.
This figure shows a photograph of an experimental setup for evaluating a user interface for directional speech extraction. The scene depicts two participants seated at opposite ends of a table in an indoor environment. Target A, wearing a light blue shirt, is positioned on the left side of the table, while Target B, wearing a green shirt, is positioned on the right side. In the foreground of the image, a smartphone-based user interface device is placed on the table, connected to what appears to be a microphone array with visible wiring. Two semi-transparent colored overlays emanate from the user interface device toward each target speaker: a cyan/blue beam extending toward Target A and a green beam extending toward Target B, visually representing the directional focus capabilities of the system. These colored beams illustrate how the system can selectively focus on individual speakers based on their spatial location. The setup demonstrates the interactive nature of the user interface, where participants can select one or both speakers for audio extraction in different room configurations and acoustic environments.
Table 2. Comparison of system performance and usability across different sector resolutions
Config Resolution Preference Accuracy Confidence (Single) Confidence (Multiple)
6 Sectors 30° 0.96 0.90 0.82 1.00
9 Sectors 20° 0.88 0.87 0.70 0.86
While our objective metrics demonstrate technical performance, understanding how users interact with directional selection in real-world scenarios is critical for practical deployment. We conduct a user study evaluating the usability of our interface across different sector resolution configurations, examining both selection accuracy and user confidence when targeting speakers in multi-person environments.
We recruit 10 participants to evaluate two configurations: our default 6-sector design (30° resolution) and a 9-sector design (20° resolution). As shown in Fig. 15, participants were positioned in three conference rooms with two experimenters as target speakers. For each configuration, participants completed three tasks: selecting target A, target B, and both targets simultaneously. Experimenter positions remained consistent across configurations while task order was randomized to minimize learning effects.
Participants were instructed to select single or multiple sectors they believed necessary to capture the target speaker(s). We evaluated interface usability through three metrics: (1) Selection Accuracy, measured as the probability that selected sectors correctly included the target speakers. (2) User Confidence, assessed on a 5-point scale (5=very confident, 1=not confident at all) for both single and multiple speaker selections. (3) Overall Preference, rated on a 5-point scale (5=strongly prefer to use, 1=strongly disprefer to use) comparing the two configurations.
Table 2 presents our results by normalizing the scores and accuracy to 0–1. The results reveal that the 6-sector configuration consistently outperforms the 9-sector design across all metrics, achieving higher selection accuracy (0.90 vs. 0.87), stronger user preference (0.96 vs. 0.88), and greater confidence for both single-sector (0.82 vs. 0.70) and multiple-sector (1.00 vs. 0.86) scenarios. These findings suggest that finer angular resolution does not necessarily improve usability—instead, the increased cognitive load of managing more sectors can reduce both accuracy and confidence.
We observed two key factors driving this preference. First, the coarser 6-sector resolution provides sufficient spatial granularity for typical meeting and lecture scenarios where speakers are distributed around a table or room, while remaining simple enough for rapid selection without requiring precise angular estimation. Second, participants reported that the 9-sector interface felt ”overwhelming” and made them second-guess their selections, particularly when attempting to capture multiple speakers. The perfect confidence score (1.00) for multiple-sector selection with 6 sectors suggests that users found it intuitive to select adjacent sectors to create broader capture regions when needed.
These results validate our design choice of 6 sectors as providing an effective balance between spatial resolution and interaction simplicity. For future work, adaptive interfaces that automatically adjust sector granularity based on speaker density could potentially combine the benefits of both approaches.
@@ 5. LIMITATIONS AND DISCUSSION
Dual-microphone design rationale. SonicSieve requires two closely spaced microphones for robust directional speech extraction. While the microstructure embeds direction-dependent spectral cues into incoming sound, these cues become entangled with the source’s intrinsic spectrum and environmental reflections in single-channel recordings. A dual-microphone design provides a ”before/after” comparison—one channel shaped by the microstructure, one unmodified—that allows the neural network to factor out source/room differences and isolate the directional signature. Our choice of an in-line microphone as the second channel offers two key advantages: (1) It can be positioned very close to the phone’s built-in top microphone (within 1 cm in our prototype), which our ablations (Fig. 8) show is critical for cross-environment robustness; (2) It is device-agnostic and can work with smartphones that have a standard audio jack or USB-C port.
@@
  • Generalization across microphone frequency responses. To enable generalization across different smartphones microphones, device-specific calibration can be performed by measuring the frequency response of the built-in and in-line microphones, and deriving equalization weights to normalize for their effects. This can allow our system to robustly perform directional speech extraction across a range of devices. A complementary approach is the use of established data-driven approaches such as training on diverse device recordings (Li et al., 2017a; Ardila et al., 2019) and targeted data augmentation techniques (Morocutti et al., 2023; Yuan et al., 2025a) which may enable zero-shot generalization without calibration for unseen devices.
@@
Pathways to device-native integration. While our current prototype uses a microstructure attached to an external wired earphones, the same core concept can be integrated directly into devices with two closely located microphones. Several integration pathways are feasible: (1) Smartphones with a top and a rear microphone near the camera (e.g., Google Pixel 7) often have <<2 cm spacing. A microstructure could be embedded near the rear mic while using the top mic as reference (Fig. 16a). (2) Standalone voice recorders with dual front-facing microphones could house the microstructure in a compact form factor (Fig. 16b). (3) Future smartphone designs could incorporate a miniaturized microstructure as part of the device chassis, with one microphone inside the structure and one outside.
Source separation within a sector. Our system currently performs spatial filtering along predefined directional sectors and produces a mixture of all sounds it. This can be undesirable when multiple speakers are close to each other and fall within the same sector. Future extensions of this work could address this limitation in two ways: (1) A low-latency speech separation framework (Gu et al., 2021, 2022; Yuan et al., 2024a) can be applied after directional speech extraction, using permutation-invariant training to separate multiple speakers. (2) More fine-grained separation can be performed in the spatial domain by building on techniques for region-customizable sound extraction (Gu and Luo, 2024) to spatially isolate the target speaker.
@@
A limitation of the current in-line microphone setup is the close proximity of the mic to the earphones, resulting in insufficient wire length to comfortably reach a user’s ears when the smartphone is placed on a table. This can be addressed in two ways. First, a custom but low-cost headphone could be designed with the in-line mic positioned farther from the speakers, allowing the phone to remain on a surface while maintaining sufficient wire length. Alternatively, wireless headphones could be used for playback by integrating the acoustic microstructure onto one of the outward-facing microphones typically used for noise cancellation. This approach is likely possible as modern headphones often include dual outward-facing microphones for noise cancellation, and low-power GPUs are becoming integrated into wearable devices (T. Tambe, E. Yang, G. G. Ko, Y. Chai, C. Hooper, M. Donato, P. N. Whatmough, A. M. Rush, D. Brooks, and G. Wei (2022); 20).
Human-centered interaction and deployment considerations. While our user study (Sec. 4.2 and 4.3) demonstrates the system performance and UI usability in controlled settings, broader human-centered aspects require deeper investigation. Real-world conversational dynamics, including natural turn-taking, situational awareness of non-target speakers, and adaptation to speaker movement, present challenges beyond spatial selection. Accessibility considerations, such as how users with hearing challenges interact with directional audio interfaces, require dedicated study. Similarly, longitudinal in-the-wild deployment across diverse acoustic environments would reveal how interaction patterns evolve over time. Our current work establishes the technical foundation and demonstrates basic usability. These human-centered dimensions represent valuable directions for future research toward broader real-world adoption.